OpenStudy (anonymous):

integral of x^4sinx/1+x^0dx in the interval of (-pi/2, pi/2)

sam (.sam.):

\[\int\limits\limits \frac{x^4 \sin (x)}{1+x^0} \, dx\] \[\int\limits\limits \frac{x^4 \sin (x)}{1+1} \, dx\] \[\int\limits\limits \frac{x^4 \sin (x)}{2} \, dx\]

sam (.sam.):

Factor \[\frac{1}{2}\int\limits x^4 \sin (x) \, dx\] Then use integration by parts 4 times

OpenStudy (anonymous):

how would you do u sub?

sam (.sam.):

u=x^4 dv=sin(x)dx du=4x^3dx v=-cos(x) \[-x^4cos(x)+\int\limits\limits_{}^{}4x^3\cos(x)~dx\]